OpenStudy (anonymous):

What are the next three terms of the sequence : -30, -14,-6,-2,0?

OpenStudy (anonymous):

it looks like you increase by an amount to get to the next term... but the amount you increase changes with each term. Can you see the pattern?

OpenStudy (anonymous):

no, I dont see it, can you explain it to me ?? please!

OpenStudy (anonymous):

almost the right idea, but I think you are going the wrong direction... what are the next three terms after 0? Should be getting larger, I think...

OpenStudy (anonymous):

-30 ---> -14 increases by 16 -14 ---> -6 increases by 8 and so on...

OpenStudy (anonymous):

Almost.. but you broke the pattern :) to get from -6 to -2, increase by 4 to get from -2 to 0, increase by 2 (half as much as last increase) So... to get from 0 to the next term, increase by ___ to get ____

OpenStudy (anonymous):

Ok, so.. I got that one...do ya'll know the answer to this one.....-4x+16/x2-9 ?

OpenStudy (anonymous):

I think @haleym96 is still working through it... why don't you post this as a new question?

OpenStudy (anonymous):

@JakeV8 So the next term is 1?

OpenStudy (anonymous):

yeah, you're getting it now :) How about the next two terms?

OpenStudy (anonymous):

.5 and .25

OpenStudy (anonymous):

well... those sound like the amounts of the increases, right? not the actual terms in the sequence? The term you just found was "1". So the next increase from 1 gives you the second term.

OpenStudy (anonymous):

So, 1.5 and 1.25.

OpenStudy (anonymous):

hehe... so so close... Remember, it's increasing every time. So the last term given was 0, and you increased correctly by 1 to get 1. Then you increased again correctly by 0.5 to get 1.5 Finally, you increase FROM 1.5 by 0.25 to get ?? :)

OpenStudy (anonymous):

Haha yeah, I got it now.

OpenStudy (anonymous):

Good for you... just have to concentrate and not lose focus as you work through the pattern. Was this homework for you too, or just something you were watching because it was interesting?

OpenStudy (anonymous):

Just something I was trying to figure out myself.

OpenStudy (anonymous):

I'm doing my homework, watching tv, and doing this so it got confusing haha.

OpenStudy (anonymous):

@Greciamf so the next three terms are 1, 1.5, and 1.75

OpenStudy (anonymous):

Awesome :) glad you stuck with it... nice to see... keep it up! :)

OpenStudy (anonymous):

-30,-14,-6,-2,0? -30&-14=16 -14&-6=8 -6&-2=4 -2&0=2 anwers: 0&1=1 1&1 1/2=1/2 1 1/2&1 3/4=1/4 bonus: 1 3/4&1 7/8=1/8 -30,-14,-6,-2,0,1,1 1/2,1 3/4,1 7/8